Mariposa, Inc., produces machine tools and current
Answer #1

total OH costs= plant A costa+ plant b costs= 720000+180000=900000

total no. of machine hours= plant a hrs + plant b hrs= 120000+60000=180000

predetermined OH rate= estimated OH costs / total no. of hours= 900000/180000=$5 per hr

predetermined OH rate for Plant A= 720000/120000= $6 per hour

predetermined Oh rate for Plant B= 180000/60000=$3 per hr

OH cost assigned to each product using plantwide OH rate:

ProductX75= (60+150)*5=1050$

Product Y15= (150+60)*5=1050$

OH computation based on department OH rate:

ProductX75= 60*6 +150*3=810

Product Y15= 150*6+60*3=1080

Assignment of department wise OH rate is better as Oh can be allocated correctly among the products based on their usage

OH cost of Department B= 360000

OH rate of Department B = 360000/60000= 6$ per hour

OH computation based on department OH rate:

ProductX75= 60*6 +150*6=1260

Product Y15= 150*6+60*6=1260

Some companies have moved beyond both the plant-wide rate and the departmental rates because they want to consider all of the activities that are driving up manufacturing overhead costs. These companies are attempting to find the root causes of the indirect manufacturing costs so they can assign the costs to products in a more logical manner (instead of merely spreading the costs arbitrarily via plant-wide or departmental production hours). This effort is known as activity based costing.

Single Overhead Rates

Single overhead rates apply cost allocations for expenses incurred across the entire plant. Common overhead rates include cost allocations for such expenses as indirect materials, indirect labor, utilities expenses, insurance and property taxes, as well as depreciation of buildings, furnishings and equipment. Single overhead rates are figured by dividing the total cost of overhead by cost drivers common throughout each department or section of the business. Cost drivers can be defined as the primary reasons for indirect costs.

Separating Rates

Departmental and manufacturing overhead rates are those calculated for each separate department. This rate is figured by dividing the total department overhead budgeted by the budgeted amount of the common cost drivers within the department. For example, in a manufacturing business, the machining department may use machine-hours to figure overhead rates when calculating job costs. Contrastingly, the shipping department may use labor hours to figure overhead rates. Other areas of a plant that produces multiple products may allocate overhead rates of either machine-hours or labor to the budgeted job costs depending on the main activity of each department.